Physics

A force of 40.0 N is applied in the positive x-direction to a block of mass 8.75 kg, at rest on a frictionless surface.


(a)

What is the block's acceleration (in m/s2)? (Indicate the direction with the sign of your answer.)

(b)

How fast is it going (in m/s) after 5.50 s?
Answer #1
F=ma 40 = 8.75 x a a = 40/8.75 a = 4.57 (b) v = a x t v = 4.57 x 5.5 v = 25.135ms^-1
answered by: Franklin77